OpenStudy (lukecrayonz):

Find the average rate of change for each function on the interval [-2,6]. f(x)=-x^4+3x

hero (hero):

So easy

OpenStudy (lukecrayonz):

Then help please, i don't know the formula.

hero (hero):

slope formula

OpenStudy (anonymous):

integrate the function from -2 to 6, then divide by 8

OpenStudy (anonymous):

oh actually i am wrong and hero is right

OpenStudy (anonymous):

\[\frac{f(6)-f(-2)}{6-(-2)}\]

hero (hero):

Wow, That rarely happens

OpenStudy (anonymous):

lol

OpenStudy (lukecrayonz):

So i solve and thats my answer? Thats it?

hero (hero):

yup

OpenStudy (anonymous):

f'(x) = -4 x^3 +3 the rate = -4 (-2)^2 +3 =-4(-8) +3 = 32+3=35

OpenStudy (anonymous):

?

hero (hero):

DHASHNI, wth

OpenStudy (lukecrayonz):

So i'd do f(-2^4+3*-2)-f(-6^4+18)/6+2?

OpenStudy (lukecrayonz):

((-2^4-6)-(-6^4+18)/8

OpenStudy (lukecrayonz):

137.75? Sounds right?